Determine the probability P(16 or more) for a binomial experiment with n = 18 trials and the success probability p = 0.9. Then find the mean, variance and standard deviation.
Answer #1

Sol: Given data : n=18 ; P=0.9 ; q l-p 1-0.9 q = 0.1 DP ( 16 or more P ( X =16) + P ( x = 17) +P(X=18) (where n=18 P(x = 16)plz support my work thank you for your considerations